OpenStudy (anonymous):

2x-x=5/x-1 when x is -3

OpenStudy (anonymous):

and need help whit his \[3\sqrt{27/64}\]

OpenStudy (anonymous):

is x -1 or just 1

OpenStudy (anonymous):

ok what about the other one

OpenStudy (anonymous):

your problem is wrong.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

wait never mind ok 2(-3)-(3)/(-3)-1 =-9/-4 yea wrong problem

OpenStudy (anonymous):

Thats what I believe too..

OpenStudy (anonymous):

\[3\sqrt{27/64}\] need help wit this one

OpenStudy (anonymous):

and need help whit 2xsq-x+5 /x-1 when x= -3

jhonyy9 (jhonyy9):

i have missed one minus correct is -6+3=5/-4 -3=5/-4 12=5 - so x=-3 not is solution

OpenStudy (anonymous):

27 is 3^3 and 64 is 4^3 so the answer is 3/4

jhonyy9 (jhonyy9):

the radical exercise result is 3/4

jhonyy9 (jhonyy9):

because 3*3*3=27 and 4*4*4=64

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

ok waht about this \[\sqrt{75}/2 +3\sqrt{2}/4\]

jhonyy9 (jhonyy9):

what sign this ?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

sq i think 75sq/2+3 2sq/4

OpenStudy (anonymous):

you with me

jhonyy9 (jhonyy9):

ok but this i have thought for wow what sign

jhonyy9 (jhonyy9):

so sqrt 75 = 5sqrt3 than there is 5sqrt3/2 +3sqrt2/4 =(10sqrt3 +3sqrt2)/4

jhonyy9 (jhonyy9):

this is right understandably ?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

75sq /2 + 3 and 2sq /4